# Collapsing regular Riemannian foliations with flat leaves

### v1

@@ 3.3. Proof of Theorem A
þ Let $M$ be a smooth manifold, and let $\mathcal{F}^{\prime}$ and $\mathcal{F}$ be smooth regular foliations of dimension $k$ on $M$. Assume that $\mathcal{F}$ and $\mathcal{F}^{\prime}$ agree on $D\subset M$ a dense subset of $M$, i.e. for $p\in D$ we have that $L_{p}=L_{p}^{\prime}$ where $L_{p}\in\mathcal{F}$ and $L_{p}^{\prime}\in\mathcal{F}^{\prime}$ are the respective leaves through $p$ of $\mathcal{F}$ and $\mathcal{F}^{\prime}$. Then $\mathcal{F}=\mathcal{F}^{\prime}$, that is the foliations agree everywhere.
@@
We recall that the foliations $\mathcal{F}$ and $\mathcal{F}^{\prime}$ induce smooth distributions $\Delta$ and $\Delta^{\prime}$ respectively, by setting $\Delta(q)=T_{q}L_{q}$ and $\Delta^{\prime}(q)=T_{q}L_{q}^{\prime}$. We point out that we can consider this distributions as continuous functions $\Delta,\,\Delta^{\prime}\colon M\to X$. By hypothesis these functions agree on a dense subset, and are continuous.
This in turn implies that for all $q\in M$ we have $T_{q}L_{q}=T_{q}L_{q}^{\prime}$. But observe that by construction we have that $L_{q}$ is an integral submanifold of $\Delta$, and $L_{q}^{\prime}$ is an integral submanifold of $\Delta^{\prime}$. By uniqueness of such integral submanifolds it follows that $L_{q}=L_{q}^{\prime}$. ∎
Proof of þA.
In the case when $m=1$ there are no such foliations: Assume that $(M,\mathcal{F},g)$ is a regular $B$-foliation on a compact simply-connected $1$-dimensional manifold. Since the leaves have positive dimension, we conclude that the foliation consists of only one leaf, i.e. =$\mathcal{F}=\{M\}$, which is a contradiction to $\mathcal{F}$ being non-trivial.
In the case of $m=2$, we have several cases based on the codimension of the foliation. In the case when the foliation has codimension 00, this implies that the foliation consists of only one leaf, i.e. =$\mathcal{F}=\{M\}$, which is a contradiction to the non-triviality of $\mathcal{F}$. The case of codimension $2$ is excluded, because this implies that a leaf is of dimension 00, contradicting the assumption that the leaves of $\mathcal{F}$ have positive dimension. In the case when $\mathcal{F}$ has codimension $1$, from [22, Theorem D] it follows that given an $A$-foliations of codimension $1$ on a $2$-dimensional simply-connected manifold $M$, then $M$ is diffeomorphic to $\mathbf{S}^{2}$ and $\mathcal{F}$ is given by a circle action. This is a foliation with singular leaves and contradicts our assumption of regularity.
From now on we consider the case $m>2$. Since the hypothesis of þB are satisfied, and thus there exists a collapsing sequence of Riemannian metrics $g_{\delta}$ with bounded curvature as defined in (3.1). Given $\varepsilon>0$, we consider the $\mathsf{N}$-structure $\widetilde{\eta}_{\varepsilon}$ given by þ2.15. By þ2.14 it follows that the $\mathsf{N}$-structure $\widetilde{\eta}_{\varepsilon}$ is given by an effective torus actions $\mu_{\varepsilon}\colon T^{k}\times M\to M$, and we denote by $T^{k}_{\varepsilon}\subset\mathrm{Diff}(M)$ the image of $T^{k}$ under the map $T^{k}\ni\xi\mapsto\mu_{\varepsilon}(\xi,\cdot)\in\mathrm{Diff}(M)$.
Moreover, by þ2.15 for each $\varepsilon>0$ we have a Riemannian metric $(M,g^{\varepsilon}_{N(\varepsilon)})$ which is $\mu_{\varepsilon}$-invariant. In the proof of þC we modified these metrics $g^{\varepsilon}_{N(\varepsilon)}$ to define in (3.3) a sequence of Riemannian metrics $\bar{g}^{\varepsilon}_{N(\varepsilon)}$, which by þ3.1 converges uniformly to the Riemannian metric $g$. We also highlight that by construction, the metrics $\bar{g}^{\varepsilon}_{N(\varepsilon)}$ are $\mu_{\varepsilon}$-invariant.
We have the following claim:
Claim 3.2.1.
þ The sequence $(M/T^{k}_{\varepsilon},d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})$ converges to $(M/\mathcal{F},d_{g}^{\ast})$ as $\varepsilon\to 0$.
Proof of þ3.2.1.
Let $\lambda\in\mathbb{R}$ be the lower sectional curvature bound of $(M,g)$. Given $\varepsilon>0$, we have by construction that $(M,g^{\varepsilon}_{N(\varepsilon)})$ has $\sec(g^{\varepsilon}_{N(\varepsilon)})\geq\lambda-\varepsilon$. Thus the space $(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$ is an Alexandrov space with $\operatorname{curv}\geq\lambda-\varepsilon$. Therefore, given $\varepsilon_{0}>0$, for $0<\varepsilon\leq\varepsilon_{0}$ we have that $(M/\mathcal{F},d_{g}^{\ast})$ and $(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$ are compact Alexandrov spaces with $\operatorname{curv}\geq\lambda-\varepsilon_{0}$. Moreover, they both have the same dimension, and by construction $(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$ converges to $(M/\mathcal{F},d^{\ast}_{g})$ in the Gromov-Hausdorff sense.
We fix $p^{\ast}\in M/\mathcal{F}$, and consider $q^{\ast}\in M/\mathcal{F}$ close enough to $p^{\ast}$ in $M/\mathcal{F}$, such that there is a unique geodesic $\gamma^{\ast}$ in $M/\mathcal{F}$ joining $p^{\ast}$ to $q^{\ast}$. By [27, Lemma 2.2] we can find geodesics $\gamma^{\ast}_{\varepsilon}$ in $(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$ starting at $p^{\ast}$ which converge uniformly to $\gamma^{\ast}$. We denote by $q_{\varepsilon}^{\ast}$ the other end point of the geodesic $\gamma^{\ast}_{\varepsilon}$. Observe that by construction $q_{\varepsilon}^{\ast}$ converge to $q^{\ast}$.
@@
We now consider lifts $q_{\varepsilon}$ and $\bar{q}_{\varepsilon}$ in $M$ of $q^{\ast}_{\varepsilon}$ for the projection map $\pi_{\varepsilon}\colon M\to M/T^{k}_{\varepsilon}$ such that $d_{g^{\varepsilon}_{N(\varepsilon)}}(p,q_{\varepsilon})=d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$ and $d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p,\bar{q}_{\varepsilon})=d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$. Thus, there exists $V_{\varepsilon}\in\nu^{g^{\varepsilon}_{N(\varepsilon)}}_{p}(M,T^{k}_{\varepsilon})$ and $\bar{V}_{\varepsilon}\in\nu^{\bar{g}^{\varepsilon}_{N(\varepsilon)}}_{p}(M,T^{k}_{\varepsilon})$ such that $\exp_{p}^{g^{\varepsilon}_{N(\varepsilon)}}(V_{\varepsilon})=q_{\varepsilon}$ and $\exp_{p}^{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(\bar{V}_{\varepsilon})=\bar{q}_{\varepsilon}$. By construction there exists $\xi_{\varepsilon}\in T^{k}_{\varepsilon}$ such that $\xi_{\varepsilon}q_{\varepsilon}=\bar{q}_{\varepsilon}$. We point out that $\xi_{\varepsilon}$ can consists of two elements $\bar{\xi}_{\varepsilon}\in(T^{k}_{\varepsilon})_{p}$ and $\tilde{\xi}_{\varepsilon}\not\in(T^{k}_{\varepsilon})_{p}$. In particular since $T^{k}$ is abelian we have that, there exists an action field $X^{\ast}_{\varepsilon}\in T_{p}T^{k}_{\varepsilon}(p)$ (determined by $\tilde{\xi}_{\varepsilon}$) such that $V_{\varepsilon}=D_{p}\bar{\xi}_{\varepsilon}(\bar{V}_{\varepsilon})+X_{\varepsilon}^{\ast}$. Since $T^{k}$ is compact, and the element $X^{\ast}_{\varepsilon}$ is determined by the exponential map of $T^{k}$, we conclude that $X^{\ast}_{\varepsilon}$ converges in $T_{p}M$, up to a subsequence to a vector $X$. Moreover, by Theorem C the tangent spaces of the $T^{k}_{\varepsilon}$-orbits converge to the tangent space of the leaves of $\mathcal{F}$, and thus we have that $X$ is tangent to the leaf $L_{p}$.
@@
We also have that $d_{g^{\varepsilon}_{N(\varepsilon)}}(p,q_{\varepsilon})=d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$, which converges to $d^{\ast}_{g}(p^{\ast},q^{\ast})$ as $\varepsilon$ goes to 00. This implies that the sequence $\{q_{\varepsilon}\}\subset M$ is eventually contained in a tubular neighborhood of $L_{p}$ with respect to $g$. In particular the sequence $\{q_{\varepsilon}\}$ is eventually bounded with respect to $g$. This implies that there exists $q_{0}\in M$ such that, up to a subsequence, $q_{\varepsilon}$ converges to $q_{0}$ with respect to $g$.
@@
Since $V$ realizes the distance between $L_{p}$ and $L_{q_{0}}$, then $V$ is $g$-orthogonal to $L_{p}$. But we also have that $V=\bar{V}+X$ with $X$ tangent to the leaf $L_{p}$, possibly $\bar{0}$. This implies that $X=0$, and thus $\bar{V}=V$.
Moreover, by construction we have that $\|\bar{V}_{\varepsilon}\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}=\|D_{p}\xi_{\varepsilon}(\bar{V}_{\varepsilon})\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ converges to $\|\bar{V}\|^{2}_{g}=\|V\|^{2}_{g}$. But since $\|\bar{V}_{\varepsilon}\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ $=d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$, we conclude that $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$ converges to $d_{g}^{\ast}(p^{\ast},q^{\ast})$.
Since $q^{\ast}$ is arbitrary, we conclude that for sufficiently small balls around $p^{\ast}$, the same holds. I.e. the function $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},-)$ approximates the function $d_{g}^{\ast}(p^{\ast},-)$. By the triangle inequality, and the fact that the points $q_{\varepsilon}^{\ast}$ converge to $q$ with respect to $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ and $d^{\ast}_{g}$ we have that the metrics $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ converge uniformly to $d^{\ast}_{g}$ over the sufficiently small balls around $p^{\ast}$ in $M/\mathcal{F}$. Since $M/\mathcal{F}$ is compact, we conclude that the metrics $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ converge uniformly to $d^{\ast}_{g}$. Thus, we conclude that $(M/T^{k}_{\varepsilon},d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})$ converges to $(M/\mathcal{F},d^{\ast}_{g})$ in the Gromov-Hausdorff sense. ∎
Since $(M,d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})$ converges uniformly to $(M,d_{g})$, we have by þ2.9 and þ2.12 that there exists a closed group $G$ acting effectively on $(M,d_{g})$ by distance preserving isometries. By the work of Myers and Steenrod [36] it follows that $G$ is a closed subgroup of $\mathrm{Isom}(M,g)$, the group of isometries of $g$.
Moreover by þ2.11 we have that $(M/T^{k}_{\varepsilon},d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}})$ converges to $(M/G,d_{g}^{\ast})$ in the Gromov-Hausdorff sense. By þ3.2.1, we have by the uniqueness of the limit that $(M/G,d_{g}^{\ast})$ is isometric to $(M/\mathcal{F},d^{\ast}_{g})$.
But from the proof of þ3.2.1 we obtain the following claim:
Claim 3.2.2.
þ Over the set of $G$-principal orbits, we have $T_{p}G(p)=T_{p}L_{p}$.
@@
Since $(M/\mathcal{F},d^{\ast}_{g})$ is an Alexandrov space with $\operatorname{curv}\geq\lambda$, then $(M/G,d^{\ast}_{g})$ is an Alexandrov space with the same lower curvature bound. Thus given $\varepsilon_{0}>0$, for $\varepsilon_{0}\geq\varepsilon>0$ we have that $(M/T^{k}_{g^{\varepsilon}_{N(\varepsilon)}})$ and $(M/G,d^{\ast}_{g})$ are Alexandrov spaces with $\operatorname{curv}\geq\lambda-\varepsilon_{0}$.
We consider $G(p)$ a principal orbit. Then there exists a sufficiently small ball around $G(p)$ in $(M/T^{k}_{g^{\varepsilon}_{N(\varepsilon)}})$ consisting of only $G$-principal orbits. Let $G(q)$ be an orbit sufficiently close to $G(p)$, and let $\gamma^{\ast}_{G}$ be the minimizing geodesic in the principal stratum of $(M/G,d^{\ast}_{g})$ joining $G(p)$ to $G(q)$. Then by [27, Lemma 2.2] there exists minimizing geodesics $\tilde{\gamma}^{\ast}_{\varepsilon}$ in the principal stratums of $(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$ converging uniformly to $\gamma^{\ast}_{G}$.
Take $\delta>0$ small enough so that the $\delta$-ball around $p$ in $(M,g)$ is consist of only principal orbits, and consider $B_{\delta}^{\perp}(p)\subset(M,g)$ the set of all points $x$ in $M$ with $d_{g}(p,x)<\delta$ such that there exist a unique minimizing geodesic $\gamma\colon[0,1]\to(M,g)$ with $\gamma(0)=p$, $\gamma(1)=x$, and $\gamma^{\prime}(0)$ perpendicular to $T_{p}G(p)$ with respect to $g$ (for example we can take $\delta$ smaller than the injectivity radius of $g$ at $p$). Observe that by the Slice Theorem for group actions and how the metric $d^{\ast}_{g}$ is defined on $M/G$, the open ball $B_{\delta}(G(p))\subset(M/G,d^{\ast}_{g})$ is isometric to $B_{\delta}^{\perp}(p)$ under the quotient map $\pi_{G}\colon(M,d_{g})\to(M/G,d_{g}^{\ast})$. Under this identification, the lift $\gamma_{G}$ in $B_{\delta}^{\perp}$ of $\gamma^{\ast}_{G}$ is completely determined by the lift $p$ of $G(p)$.
We also point out that since $(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$ converges in the Gromov-Hausdorff sense to $(M/G,d^{\ast}_{g})$, by þ2.7 we may assume that $p$ is contained in a $T^{k}_{\varepsilon}$-principal orbit. Thus via the quotient map $\pi_{\varepsilon}\colon(M,d_{g^{\varepsilon}_{N(\varepsilon)}})\to(M/T^{k}_{\varepsilon},d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}})$, the open ball $B^{\varepsilon}_{\delta}(p^{\ast})$ is isometric to $B_{\delta}^{\varepsilon\,\perp}(p)\subset(M,d_{g^{\varepsilon}_{N(\varepsilon)}})$, the collection of all points in $B^{\varepsilon}_{\delta}(p)\subset M$ such that there exist a minimizing geodesic $\alpha_{\varepsilon}\colon[0,1]\to(M,g)$ starting at $p$, and $\alpha_{\varepsilon}^{\prime}(0)$ perpendicular to $T_{p}T^{k}_{\varepsilon}(p)$ with respect to $g^{\varepsilon}_{N(\varepsilon)}$. We also point out that the lift $\tilde{\gamma}\varepsilon$ to $B_{\delta,\varepsilon}^{\perp}$ of $\tilde{\gamma}^{\ast}_{\varepsilon}$ is completely determined by the lift $p$ of $p^{\ast}$.
Thus we have that for each geodesic $\gamma_{G}\colon[0,1]\to M$ contained in $B_{\delta}^{\perp}(p)\subset(M,d_{g})$ there exist geodesics $\tilde{\gamma}_{\varepsilon}$ contained in $B_{\delta}^{\varepsilon\,\perp}(p)\subset(M,d_{g^{\varepsilon}_{N(\varepsilon)}})$ converging uniformly to $\gamma_{G}$.
As stated in the proof of Claim 3.2.1, the sequence of vectors $\gamma_{\varepsilon}^{\prime}(0)$ converges to a vector $V$ that is perpendicular to $L_{p}$ with respect to $g$. But $V$ is the tangent vector of a lift of the geodesic $\gamma^{\ast}$ starting at $p$. Thus we have that $\gamma^{\prime}(0)$ is a scalar multiple of $V$ and thus perpendicular to $L_{p}$. In other words, we have that $\nu^{g}_{p}(M,G(p))\subset\nu^{g}_{p}(M,L_{p})$.
@@
Since $\mathcal{F}$ is a regular Riemannian foliation we have that $M/\mathcal{F}$ is an orbifold. Given that $M/G$ is isometric to $M/\mathcal{F}$, we conclude that $M/G$ is an orbifold. Now by the Slice Theorem (see [1, Theorem 3.57]), a sufficiently small ball of $G(x)\in M/G$ in $(M/G,d^{\ast}_{G})$ is homeomorphic to $\mathbb{R}^{m}/G_{p}$. From this and the fact that $M/G$ is an orbifold, we conclude that $G_{p}$ is a finite subgroup. This implies that the Riemannian foliation on $M$ induced by the partition into the $G$-orbits is a regular Riemannian foliation. Moreover by þ3.2.2 the distributions $T_{p}L_{p}$ and $T_{p}G(p)$ agree over the open dense set of principal $G$-orbits. Then by þ3.2 we have $G(x)=L_{x}$ for an arbitrary $x\in M$.
To finish the proof we consider $x\in M$ fixed. Since we have that $G(x)$ is homeomorphic to $G/G_{x}$, and $G_{x}$ is finite, then $G$ is a covering space of $G(x)$. But $G(x)$ is also homeomorphic $L_{x}$, which is a torus. This implies that $G$ is a finite cover of a torus and thus $G$ it is aspherical with fundamental group isomorphic to $\pi_{1}(L_{x})$. Thus $G$ is a torus. ∎

### v2

@@ 3.3. Proof of Theorem LABEL:MT:_flat_foliations_are_given_by_group_actions
[…]
@@
[…]
@@
We now consider lifts $q_{\varepsilon}$ and $\bar{q}_{\varepsilon}$ in $M$ of $q^{\ast}_{\varepsilon}$ for the projection map $\pi_{\varepsilon}\colon M\to M/T^{k}_{\varepsilon}$ such that $d_{g^{\varepsilon}_{N(\varepsilon)}}(p,q_{\varepsilon})=d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$ and $d_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p,\bar{q}_{\varepsilon})=d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$. Thus, there exists $V_{\varepsilon}\in\nu^{g^{\varepsilon}_{N(\varepsilon)}}_{p}(M,T^{k}_{\varepsilon})$ and $\bar{V}_{\varepsilon}\in\nu^{\bar{g}^{\varepsilon}_{N(\varepsilon)}}_{p}(M,T^{k}_{\varepsilon})$ such that $\exp_{p}^{g^{\varepsilon}_{N(\varepsilon)}}(V_{\varepsilon})=q_{\varepsilon}$ and $\exp_{p}^{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(\bar{V}_{\varepsilon})=\bar{q}_{\varepsilon}$. By construction there exists $\xi_{\varepsilon}\in T^{k}_{\varepsilon}$ such that $\xi_{\varepsilon}q_{\varepsilon}=\bar{q}_{\varepsilon}$. We point out that $\xi_{\varepsilon}$ can be factored as $\xi_{\varepsilon}=\bar{\xi}_{\varepsilon}\cdot\tilde{\xi}_{\varepsilon}$, where $\bar{\xi}_{\varepsilon}\in(T^{k}_{\varepsilon})_{p}$ and $\tilde{\xi}_{\varepsilon}\not\in(T^{k}_{\varepsilon})_{p}$. In particular since $T^{k}$ is abelian, we have that there exists an action field $X^{\ast}_{\varepsilon}\in T_{p}T^{k}_{\varepsilon}(p)$ (determined by $\tilde{\xi}_{\varepsilon}$) such that $V_{\varepsilon}=D_{p}\bar{\xi}_{\varepsilon}(\bar{V}_{\varepsilon})+X_{\varepsilon}^{\ast}$. Since $T^{k}$ is compact, and the element $X^{\ast}_{\varepsilon}$ is determined by the exponential map of $T^{k}$, we conclude that $X^{\ast}_{\varepsilon}$ converges in $T_{p}M$, up to a subsequence, to a vector $X$. Moreover, by Theorem LABEL:MT:_Approximating_flat_foliations_by_N-structures the tangent spaces of the $T^{k}_{\varepsilon}$-orbits converge to the tangent space of the leaves of $\mathcal{F}$, and thus we have that $X$ is tangent to the leaf $L_{p}$.
@@
We also have that $d_{g^{\varepsilon}_{N(\varepsilon)}}(p,q_{\varepsilon})=d^{\ast}_{g^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$, which converges to $d^{\ast}_{g}(p^{\ast},q^{\ast})$ as $\varepsilon$ goes to 0. This implies that the sequence $\{q_{\varepsilon}\}\subset M$ is eventually contained in a tubular neighborhood of $L_{p}$ with respect to $g$. In particular the sequence $\{q_{\varepsilon}\}$ is eventually bounded with respect to $g$. This implies that there exists $q_{0}\in M$ such that, up to a subsequence, $q_{\varepsilon}$ converges to $q_{0}$ with respect to $g$.
@@
Since $V$ realizes the distance between $L_{p}$ and $L_{q_{0}}$, then $V$ is $g$-orthogonal to $L_{p}$. But we also have that $V=\bar{V}+X$ with $X$ tangent to the leaf $L_{p}$, possibly equal to $\bar{0}$. This implies that $X=\bar{0}$, and thus $\bar{V}=V$.
Moreover, by construction we have that $\|\bar{V}_{\varepsilon}\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}=\|D_{p}\xi_{\varepsilon}(\bar{V}_{\varepsilon})\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ converges to $\|\bar{V}\|^{2}_{g}=\|V\|^{2}_{g}$. But since $\|\bar{V}_{\varepsilon}\|^{2}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}$ $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$, we conclude that $d^{\ast}_{\bar{g}^{\varepsilon}_{N(\varepsilon)}}(p^{\ast},q^{\ast}_{\varepsilon})$ converges to $d_{g}^{\ast}(p^{\ast},q^{\ast})$.
[…]
We now use results from Section 2.6 to proof the following lemma.
Lemma 3.5.
There exists a Lie group $G$ acting by isometries on $(M,g)$ such that $(M/G,d^{\ast}_{g})$ is isometric to $(M/\mathcal{F},d^{\ast}_{g})$.
@@
[…]
For the group action from ?THM? LABEL:L:_existence_of_limit_group_action_for_torus_actions we obtain the following conclusion.
Lemma 3.6.
Over the set of $G$-principal orbits, we have $T_{p}G(p)=T_{p}L_{p}$.
Proof.
[…]
Take $\delta>0$ small enough so that the $\delta$-ball around $p$ in $(M,g)$ consists of only principal orbits, and consider $B_{\delta}^{\perp}(p)\subset(M,g)$ the set of all points $x$ in $M$ with $d_{g}(p,x)<\delta$ such that there exists a unique minimizing geodesic $\gamma\colon[0,1]\to(M,g)$ with $\gamma(0)=p$, $\gamma(1)=x$, and $\gamma^{\prime}(0)$ perpendicular to $T_{p}G(p)$ with respect to $g$ (for example we can take $\delta$ smaller than the injectivity radius of $g$ at $p$). Observe that by the Slice Theorem for group actions and how the metric $d^{\ast}_{g}$ is defined on $M/G$, the open ball $B_{\delta}(G(p))\subset(M/G,d^{\ast}_{g})$ is isometric to $B_{\delta}^{\perp}(p)$ under the quotient map $\pi_{G}\colon(M,d_{g})\to(M/G,d_{g}^{\ast})$. Under this identification, the lift $\gamma_{G}$ in $B_{\delta}^{\perp}$ of $\gamma^{\ast}_{G}$ is completely determined by the lift $p$ of $G(p)$.
[…]
As stated in the proof of ?THM? LABEL:Cl:_Convergence_of_modified_orbit_spaces_converges_to_leaf_space, the sequence of vectors $\gamma_{\varepsilon}^{\prime}(0)$ converges to a vector $V$ that is perpendicular to $L_{p}$ with respect to $g$. But $V$ is the tangent vector of a lift of the geodesic $\gamma^{\ast}$ starting at $p$. Thus we have that $\gamma^{\prime}(0)$ is a scalar multiple of $V$ and thus perpendicular to $L_{p}$. In other words, we have that $\nu^{g}_{p}(M,G(p))\subset\nu^{g}_{p}(M,L_{p})$.
@@
We now prove the following auxiliary lemma.
Lemma 3.7.
Let $M$ be a smooth manifold, and let $\mathcal{F}^{\prime}$ and $\mathcal{F}$ be smooth regular foliations of dimension $k$ on $M$. Assume that $\mathcal{F}$ and $\mathcal{F}^{\prime}$ agree on $D\subset M$ a dense subset of $M$, i.e. for $p\in D$ we have that $L_{p}=L_{p}^{\prime}$ where $L_{p}\in\mathcal{F}$ and $L_{p}^{\prime}\in\mathcal{F}^{\prime}$ are the respective leaves through $p$ of $\mathcal{F}$ and $\mathcal{F}^{\prime}$. Then $\mathcal{F}=\mathcal{F}^{\prime}$, that is the foliations agree everywhere.
Proof.
We recall that the foliations $\mathcal{F}$ and $\mathcal{F}^{\prime}$ induce smooth distributions $\Delta$ and $\Delta^{\prime}$ respectively, by setting $\Delta(q)=T_{q}L_{q}$ and $\Delta^{\prime}(q)=T_{q}L_{q}^{\prime}$. We point out that we can consider these distributions as continuous functions $\Delta,\,\Delta^{\prime}\colon M\to\mathrm{Gr}(k,m)$. By hypothesis these functions agree on a dense subset, and are continuous.
This in turn implies that for all $q\in M$ we have $T_{q}L_{q}=T_{q}L_{q}^{\prime}$. But observe that by construction we have that $L_{q}$ is an integral submanifold of $\Delta$, and $L_{q}^{\prime}$ is an integral submanifold of $\Delta^{\prime}$. By the uniqueness of such integral submanifolds it follows that $L_{q}=L_{q}^{\prime}$. ∎
Lemma 3.8.
For the group action of $G$ on $(M,g)$ given by ?THM? LABEL:L:_existence_of_limit_group_action_for_torus_actions, we have that $L_{x}=G(x)$ and $G$ is a torus.
Proof.
We begin by pointing out that since $\mathcal{F}$ is a regular Riemannian foliation, we have that $M/\mathcal{F}$ is an orbifold. Given that $M/G$ is isometric to $M/\mathcal{F}$ by ?THM? LABEL:L:_existence_of_limit_group_action_for_torus_actions, we conclude that $M/G$ is an orbifold. Now by the Slice Theorem (see [1, Theorem 3.57]), a sufficiently small ball of $G(x)\in M/G$ in $(M/G,d^{\ast}_{G})$ is homeomorphic to $\mathbb{R}^{m}/G_{p}$. From this and the fact that $M/G$ is an orbifold, we conclude that $G_{p}$ is a finite subgroup. This implies that the Riemannian foliation on $M$ induced by the partition into the $G$-orbits is a regular Riemannian foliation. Moreover by ?THM? LABEL:Cl:_Tangent_spaces_of_G_principal_orbits_agree_with_tangent_spaces_of_leaves the distributions $T_{p}L_{p}$ and $T_{p}G(p)$ agree over the open dense set of principal $G$-orbits. Then by ?THM? LABEL:L:_singular_Riemannian_foliation_determined_on_principal_part we have $G(x)=L_{x}$ for an arbitrary $x\in M$.
To finish the proof we consider a fixed $x\in M$. Since we have that $G(x)$ is homeomorphic to $G/G_{x}$, and $G_{x}$ is finite, then $G$ is a covering space of $G(x)$. But $G(x)$ is also homeomorphic to $L_{x}$, which is a torus. This implies that $G$ is a finite cover of a torus and thus $G$ it is aspherical with fundamental group isomorphic to $\pi_{1}(L_{x})$. Thus $G$ is a torus. ∎
From the discussion above, we have proven the following lemma.
Lemma 3.9.
Let $(M,\mathcal{F},g)$ be a regular $B$-foliation on a compact-simply connected Riemannian manifold of dimension >2>2. Then $\mathcal{F}$ is given by a torus action by isometries on $(M,g)$.
Proof of ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions.
By ?THM? LABEL:L:_case_of_dimension_>_2 we have proven ?THM? LABEL:MT:_flat_foliations_are_given_by_group_actions in the case when $m>2$. We now prove the missing lower dimensional cases.
In the case when $m=1$ there are no such foliations: Assume that $(M,\mathcal{F},g)$ is a regular $B$-foliation on a compact simply-connected $1$-dimensional manifold. Since the leaves have positive dimension, we conclude that the foliation consists of only one leaf, i.e. $\mathcal{F}=\{M\}$, which is a contradiction to $\mathcal{F}$ being non-trivial.
In the case of $m=2$, we have several cases based on the codimension of the foliation. In the case when the foliation has codimension 0, this implies that the foliation consists of only one leaf, i.e. $\mathcal{F}=\{M\}$, which is a contradiction to the non-triviality of $\mathcal{F}$. The case of codimension $2$ is excluded, because this implies that a leaf has dimension 0, contradicting the assumption that the leaves of $\mathcal{F}$ have positive dimension. In the case when $\mathcal{F}$ has codimension $1$, from [22, Theorem D] it follows that given an $A$-foliation of codimension $1$ on a $2$-dimensional simply-connected manifold $M$, then $M$ is diffeomorphic to $\mathbf{S}^{2}$ and $\mathcal{F}$ is given by a circle action. This is a foliation with singular leaves and contradicts our assumption of regularity. ∎